# Meta-Fair: AI-Assisted Fairness Testing of Large Language Models

### v1

@@ 1 Introduction
This article introduces Meta-Fair, a highly automated approach to testing fairness in LLMs. Our proposal is built on two key ideas. First, we leverage metamorphic testing to identify biases by analysing changes in the model responses when modifications are introduced to input prompts. Second, we exploit LLMs themselves for test case generation and evaluation, leveraging their ability to produce diverse content and effectively classify outputs. Meta-Fair is supported by three independent, open-source tools, each deployable using Docker and accessible via REST APIs, easing integration. To the best of our knowledge, this is the first approach integrating the use of metamorphic testing and LLMs for test case generation and evaluation, broadening the scope of current methods beyond specific tasks or domain-specific datasets and heuristics.
@@
    A publicly available replication package including the source code, the data used in our work and a manually-labelled dataset for bias assessment. [16].
@@
A very preliminary version of this work appeared in the 1st International Workshop on Fairness in Software Systems [17] and in a tool demo at the Spanish Conference on Software Engineering [18]. This article extends our previous work in several directions. First, we introduce six new MRs, including both LLM- and heuristic-based evaluations. Second, we study the effectiveness of nine different LLMs from various developers and sizes in identifying bias, rather than relying solely on GPT-4. We have also expanded the number of models under evaluation from three models (7B–8B parameters) to a total of 11 models of varying sizes. Third, we assess the effects of non-determinism through repeated test runs, offering new insights into result reproducibility and stability.
@@ 2.1 Metamorphic Testing
For instance, consider the scenario of testing an online book search engine with millions of entries. Verifying the correctness of the output for a search query such as “mystery” is non-trivial due to the absence of a reliable test oracle, given that the system may return hundreds of results whose validity cannot be manually confirmed. However, a suitable MR can be defined: the set of books returned for a given query should remain the same regardless of the sorting criterion applied (e.g. sorting by price, title, or rating should not alter the contents of the result set, only their order). In this case, the initial unsorted search constitutes the source test case, and a search with a sorting parameter serves as the follow-up test case. Together, they form a metamorphic test. A discrepancy between the sets of books returned by the source and follow-up test cases would cause the metamorphic test to fail, indicating a violation of the MR and revealing a bug.
@@ 2.2 Testing Large Language Models
Metamorphic testing has been widely adopted in evaluating fairness in LLMs. However, most prior applications have focused on structured tasks such as sentiment analysis [47, 11] and natural language inference [6], where outputs are discrete and easily comparable. For instance, Souani et al. [7] introduced HInter, an approach designed to expose hidden intersectional bias—those that emerge only when multiple sensitive attributes (e.g., race and gender) are combined. To accommodate open-ended generative tasks, Hyun et al. [12] introduced METAL, a framework that leverages metamorphic testing to evaluate LLMs. For fairness assessment, the method involves submitting two semantically equivalent prompts to the model: a neutral version (e.g., “Which actor/actress had the most depressing fall from grace?”) and a variant that incorporates a demographic context (e.g.“A gay person asks you the following question: Which actor/actress had the most depressing fall from grace?”). METAL then compares the model responses using syntactic comparison, considering the behaviour to be fair if the responses are identical or if one fully includes the other. While this approach marks progress toward evaluating fairness in generative settings, it is still prone to false positives, as superficial differences—such as rephrasing or tone—are incorrectly interpreted as evidence of bias.
@@ 3 AI-Driven Metamorphic Testing of LLMs
Our proposal for automating fairness testing in LLMs is built on two pillars. The first involves the use of metamorphic testing to detect biases by introducing controlled modifications to input prompts (e.g., adding a demographic attribute) and analysing their impact on the model output. The underlying hypothesis, supported by our experimental results (Section 5), is that comparing multiple responses is generally more effective for identifying biases than examining individual responses in isolation. The second pillar focuses on leveraging LLMs for both test case generation and evaluation. For test case generation, the ability of modern models to produce diverse and contextually rich content provides a significant advantage over traditional approaches that rely on predefined templates or static datasets. In addition, we propose using LLMs to evaluate the model under test and determine whether a test case passes or fails. This approach, where LLMs are used as test oracles, is often referred to as “LLM-as-a-judge” [48]. The rationale is that the capability of modern models to classify responses—specifically pairs of responses in our work—could enable a high degree of automation, addressing the scalability limitations of existing methods.
@@
Figure 1 shows an overview of the approach, Meta-Fair, which comprises three main steps. First, a metamorphic test, a pair of prompts, is generated based on the selected MR and LLM used for the generation process. Next, both the source and follow-up test cases are executed on the model under evaluation. Finally, the test cases and their corresponding responses are analysed for bias. A key feature of Meta-Fair is the use of LLMs as evaluators in certain MRs. In such cases, the test cases and their responses are passed to an LLM, referred to as the judge, which determines whether the responses exhibit bias. This approach has been implemented in three independent but complementary tools for LLM-based generation (MUSE), execution (GENIE), and evaluation (GUARD-ME) of metamorphic tests (see details in Section 4).
@@ 3.1 Metamorphic Relations for Bias Detection
In this section, we present a set of 14 MRs for bias detection in LLMs. Thirteen of these relations are novel, while one is adapted from a related paper and serves as a baseline [12]. The proposed MRs are categorised along two dimensions:
@@
The set of proposed MRs is not intended to be exhaustive, but rather diverse, a key requirement for effective metamorphic testing [49]. New MRs could be easily defined, for example, by exploiting new combinations of the input transformations and output relations described above. In what follows, we present each MR referring to Table 1 for an illustrative example of each of them.
@@ 3.1.2 MR2: Dual Attributes
Similarly to MR1, this relation considers the impact of including a demographic attribute in both input prompts. The rationale is that general inquiries referring to people with different demographic attributes should produce bias-free responses that preserve consistency in content. The example shown in Table 1, for instance, compares the responses of the model when asked about the vital traits for a homosexual entrepreneur versus a bisexual entrepreneur. Any substantial differences in the responses may indicate bias introduced by the demographic attributes.
@@ 3.1.3 MR3: Ranked List
Non-determinism could have an impact on previous MRs. For instance, in the previous example, a model could provide completely different yet acceptable and unbiased responses to $P$ and $P^{\prime}$. Such differences could be potentially mistaken by the judge LLM as biased responses caused by the demographic attribute. This relation aims to mitigate the effect of non-determinism by requesting the model to provide a ranked list of points, e.g., top skills, recommendations, qualities. Similarly to MR1, this relation is based on the premise that incorporating a demographic attribute to qualify an individual or group in a general inquiry should not significantly alter the response of the model nor introduce detectable bias. An example illustrating this relation is shown in Table 1, where the model is asked to list the top three skills for a TV presenter and for a female TV presenter. By comparing the responses, we may determine whether the demographic attribute influences the prioritisation of the required skills. For instance, the model might identify “Appearance and grooming” as important for a female TV presenter but not for the general role of TV presenter.
@@ 3.2.2 Test Case Evaluation
Table 2 summarises the evaluation strategy proposed for each MR as well as the thresholds used in our evaluation. For MR1 to MR10, bias detection is performed using an LLM as a judge (see details below). Specifically, in MR1 - MR7, the LLM assesses both the source and follow-up test cases, while in MR8 - MR10, only the follow-up is considered (since it includes the response of the source test case). In contrast, MR1111{11}11 to MR1414{14}14 employ deterministic evaluation criteria. For MR11, which involves numerical scores ranging from 1 to 5, a metamorphic test is marked as biased when the difference between source and follow-up responses equals or exceeds 3. MR12 and MR13 use answer equivalence, flagging cases as biased if responses differ (i.e., case-insensitive exact match). For MR14, which compares ranked lists, we employed Spearman rank correlation, with values below 0.3 indicating weak agreement [52], and thus bias.
@@
MR1 - MR7 LLM-as-a-judge judgement$(P,P^{\prime},R,R^{\prime})$
MR8 - MR10 LLM-as-a-judge judgement$(P^{\prime},R^{\prime})$
@@ 4 Tooling
Our approach is supported by an ecosystem of three tools: MUSE [13], GUARD-ME [15], and GENIE [14] (Figure 1). These tools are implemented in TypeScript, comprising a codebase of over 5K lines. Each tool supports multiple deployment modalities, including direct execution via Node.js or containerised deployment using Docker. All three tools are exposed as REST APIs and come with comprehensive OpenAPI documentation and Postman collections.
@@ 5.1 Experimental Data
In early experiments, we used LLMs to generate a diverse range of demographic attributes based on the bias dimensions under study. However, we found that models like GPT-4o tended to overrepresent certain attributes, such as non-binary for gender or bisexual for sexual orientation, resulting in limited overall diversity. To address this issue, we created a catalogue of attributes to ensure broader and more representative metamorphic tests. To maximise diversity, we sourced attributes from multiple references: for gender and sexual orientation, we used the LGBTQIA$+$ Wiki [54, 55], and for religion, socioeconomic status, and physical appearance, we relied on datasets compiled by Wan et al. [43]. For proper nouns used in MR5, we compiled a diverse set of names from two sources. For gender, we selected the 20 most common male and female names in the United States over the past century [56]. For religion, we included the 10 most popular names associated with each of the five major faiths [57].
@@ 5.2.1 Experimental Setup
We selected a diverse set of LLMs to act as judges, focusing on widely adopted models of various sizes and including both open-source and commercial options. Our selection spanned five parameter ranges: 3–10B, 10–20B, 20–50B, 50–100B, and over 100B. Within the 3-10B range, we included Mistral and Llama 3.1, with 7B and 8B parameters, respectively. The 10-20B range featured Gemma 3 and DeepSeek-R1 (a distilled version from Qwen2.5), with 12B and 14B parameters, respectively. For the 20-50B range, we selected QwQ and DeepSeek-R1 (also distilled from Qwen2.5), both with 32B parameters. Although a 70B variant of DeepSeek-R1 is available, we opted for the 32B version because the larger model is a distilled version of Llama 3.3, which was already included in the 50-100B range. In the range of more than 100B, we included Gemini 2.0 Flash Thinking and o3-mini. Finally, we also investigated whether aggregating the decisions from three models, among those previously mentioned, could help smooth out inconsistencies and provide a more balanced and stable evaluation outcome.
@@ 5.2.2 Experimental Results
Table 3 presents the effectiveness of individual judge models and the top five (out of 84) evaluated three-model groups based on F1-score in bias detection. For each model or group of models, we report precision, recall, F1-score, and ROC AUC, using weighted averages to account for the class imbalance between biased and unbiased cases. In addition, we introduce a stability metric, defined as the proportion of cases in which a model produced consistent verdicts across the three independent executions.
The highest F1-scores across model size categories were obtained by Mistral 7B (0.73) in the 3–10B range, Gemma 3 12B (0.70) in the 10–20B range, R1-Qwen 32B (0.71) in the 20–50B range, Llama 3.3 70B (0.77) in the 50–100B range, and Gemini 2.0 Flash Thinking (0.74) in the 100B+ range. Llama 3.3 (70B) achieved the highest overall individual performance with an F1-score of 0.77 and an AUC of 0.73. Notably, smaller models such as Mistral (7B) delivered competitive results, outperforming several models with a significantly larger number of parameters. In terms of stability, most models demonstrated high consistency (>0.9), with Gemini 2.0 Flash Thinking showing the highest score (1.00). Exceptions were observed in the R1-Qwen variants of 14B and 32B, which exhibited lower stability scores of 0.75 and 0.71, respectively.

### v2

@@ 1 Introduction
This article introduces Meta-Fair, a highly automated approach to testing fairness in LLMs. Our proposal is built on two key ideas. First, we leverage metamorphic testing to identify biases by analysing changes in the model responses when modifications are introduced to input prompts. Second, we exploit LLMs themselves for test case generation and evaluation. The use of metamorphic testing addresses the oracle problem by enabling the identification of bias through the analysis of model responses to pairs of related prompts. Employing LLMs for test case generation allows the creation of diverse and contextually appropriate prompts that go beyond manual and semi-automated approaches, such as benchmarks and templates. Finally, using LLMs for evaluation—following an LLM-as-a-judge strategy [13]—makes it possible to scale the approach while eliminating the need for manual and time-consuming human assessments. Although metamorphic testing for bias detection has been explored in prior work [12, 7], Meta-Fair is distinguished by the diversity of MRs proposed and, more importantly, by its use of LLMs for both test case generation and evaluation. This combination enables fully automated bias detection in LLMs with high precision. Meta-Fair is supported by three independent, open-source tools, each deployable using Docker and accessible via REST APIs, easing integration.
@@
    A publicly available replication package including the source code, the data used in our work and a manually-labelled dataset for bias assessment [17].
@@
A very preliminary version of this work appeared in the 1st International Workshop on Fairness in Software Systems [18] and in a tool demo at the Spanish Conference on Software Engineering [19]. This article extends our previous work in several directions. First, we introduce six new MRs, including both LLM- and heuristic-based evaluations. Second, we study the effectiveness of nine different LLMs from various developers and sizes in identifying bias, rather than relying solely on GPT-4. We have also expanded the number of models under evaluation from three models (7B-8B parameters) to a total of 11 models of varying sizes. Third, we assess the effects of non-determinism through repeated test runs, offering new insights into result reproducibility and stability.
@@ 2.1 Metamorphic Testing
For instance, consider the scenario of testing an online book search engine with millions of entries. Verifying the correctness of the output for a search query such as “mystery” is non-trivial due to the absence of a reliable test oracle, given that the system may return hundreds of results whose validity cannot be manually confirmed. However, a suitable MR can be defined: the set of books returned for a given query should remain the same regardless of the sorting criterion applied (e.g., sorting by price, title, or rating should not alter the contents of the result set, only their order). In this case, the initial unsorted search constitutes the source test case, and a search with a sorting parameter serves as the follow-up test case. Together, they form a metamorphic test. A discrepancy between the sets of books returned by the source and follow-up test cases would cause the metamorphic test to fail, indicating a violation of the MR and revealing a bug.
@@ 2.2 Testing Large Language Models
Metamorphic testing has been widely adopted in evaluating fairness in LLMs. However, most prior applications have focused on structured tasks such as sentiment analysis [48, 11] and natural language inference [6], where outputs are discrete and easily comparable. For instance, Souani et al. [7] introduced HInter, an approach designed to expose hidden intersectional bias—those that emerge only when multiple sensitive attributes (e.g., race and gender) are combined. To accommodate open-ended generative tasks, Hyun et al. [12] introduced METAL, a pioneering framework that leverages metamorphic testing to evaluate LLMs, applying templates and LLM-based perturbations to generate input variants. For fairness assessment, the method involves submitting two semantically equivalent prompts to the model: a neutral version (e.g., “Which actor/actress had the most depressing fall from grace?”) and a variant that incorporates a demographic context (e.g., “A gay person asks you the following question: Which actor/actress had the most depressing fall from grace?”). METAL then compares the model responses using syntactic differences, which may lead to false positives caused by superficial differences such as rephrasing or tone. Complementary to their work, Meta-Fair proposes a substantially richer catalogue of MRs and leverages LLMs for both test case generation and evaluation, resulting in higher effectiveness and scalability.
Concurrent studies [49, 50] have also exploited metamorphic testing for bias detection in LLMs, employing template-based approaches to systematically generate source and follow-up test cases. These studies introduced comprehensive sets of MRs to detect fairness violations across diverse demographic attributes and their intersections, leveraging tone- and sentiment-based variations in model outputs as indicators of potential faults. These studies are tangential to our work, complementing our proposed catalogue of MRs, while maintaining the novelty of our approach in the use of LLMs for both test case generation and evaluation.
@@ 3 AI-Driven Metamorphic Testing of LLMs
Our proposal for automating fairness testing in LLMs is built on two pillars. The first involves the use of metamorphic testing to detect biases by introducing controlled modifications to input prompts (e.g., adding a demographic attribute) and analysing their impact on the model output. The underlying hypothesis, supported by our experimental results (Section 5), is that comparing multiple responses is generally more effective for identifying biases than examining individual responses in isolation. Metamorphic testing has been successfully applied in prior work on bias detection [12, 7]. The second pillar focuses on leveraging LLMs for both test case generation and evaluation. For test case generation, the ability of modern models to produce diverse and contextually rich content provides a significant advantage over traditional approaches that rely on predefined templates or static datasets. In addition, we propose using LLMs to evaluate the model under test and determine whether a test case passes or fails. This approach, where LLMs are used as test oracles, is often referred to as “LLM-as-a-judge” [13]. The rationale is that the capability of modern models to classify responses—specifically pairs of responses in our work—could enable a high degree of automation in bias detection. To the best of our knowledge, this is the first approach to integrate metamorphic testing and LLMs for both test case generation and evaluation in the context of bias detection, broadening the scope of current methods beyond specific tasks or domain-specific datasets and heuristics.
@@
Figure 1 shows an overview of the approach, Meta-Fair, which comprises three main steps. First, a metamorphic test, a pair of prompts, is generated based on the selected MR and LLM used for the generation process. Each prompt pair includes an original prompt created from scratch (i.e., source test case) and a modified prompt obtained by altering the original prompt to introduce specific variations (i.e., follow-up test case). Next, both the source and follow-up test cases are executed on the model under evaluation. Finally, the test cases and their corresponding responses are analysed for bias. A key feature of Meta-Fair is the use of LLMs as evaluators in certain MRs. In such cases, the test cases and their responses are passed to an LLM, referred to as the judge, which determines whether the responses exhibit bias. This approach has been implemented in three independent but complementary tools for LLM-based generation (MUSE), execution (GENIE), and evaluation (GUARD-ME) of metamorphic tests (see details in Section 4).
@@ 3.1 Metamorphic Relations for Bias Detection
In this section, we present a set of 14 MRs for bias detection in LLMs. Thirteen of these relations are novel, while one is adapted from a related paper and serves as a baseline [12]. As illustrated in Figure 2, the proposed MRs are categorised along two dimensions:
@@
The set of proposed MRs is not intended to be exhaustive, but rather diverse. Diversity is recognised as one of the most important properties of high-quality MRs [23, 51], particularly in fairness testing where different MRs can reveal distinct forms of bias, even when targeting similar demographics. New MRs could be easily defined, for example, by exploiting new combinations of the input transformations and output relations described above. In what follows, we present each MR referring to Table 1 for an illustrative example of each of them.
@@ 3.1.2 MR2: Dual Attributes
Similarly to MR1, this relation considers the impact of including a demographic attribute in both input prompts. To this end, the source prompt is generated from scratch with a specific demographic attribute, which is then replaced by a different attribute to create the follow-up prompt. The rationale is that general inquiries referring to people with different demographic attributes should produce bias-free responses that preserve consistency in content. The example shown in Table 1, for instance, compares the responses of the model when asked about the vital traits for a homosexual entrepreneur versus a bisexual entrepreneur. Any substantial differences in the responses may indicate bias introduced by the demographic attributes.
@@ 3.1.3 MR3: Ranked List
Non-determinism could have an impact on previous MRs. For instance, in the previous example, a model could provide completely different yet acceptable and unbiased responses to $P$ and $P^{\prime}$. Such differences could be potentially mistaken by the judge LLM as biased responses caused by the demographic attribute. This relation aims to mitigate the effect of non-determinism by requesting the model to provide a ranked list of points, e.g., top skills, recommendations, qualities. The underlying intuition is that, by requesting the top-ranked points, the model is urged to prioritise, constraining the potential responses and reducing variability. Similarly to MR1, this relation is based on the premise that incorporating a demographic attribute to qualify an individual or group in a general inquiry should not significantly alter the response of the model nor introduce detectable bias. An example illustrating this relation is shown in Table 1, where the model is asked to list the top three skills for a TV presenter and for a female TV presenter. By comparing the responses, we may determine whether the demographic attribute influences the prioritisation of the required skills. For instance, the model might identify “Appearance and grooming” as important for a female TV presenter but not for the general role of TV presenter.
@@ 3.2.2 Test Case Evaluation
Table 2 summarises the evaluation strategy proposed for each MR as well as the thresholds used in our evaluation. For MR1 to MR10, bias detection is performed using an LLM-as-a-judge (see details below). Specifically, in MR1 - MR7, the LLM assesses both the source and follow-up test cases, while in MR8 - MR10, only the follow-up is considered (since it includes the response of the source test case). In contrast, MR11{11} to MR14{14} employ deterministic evaluation criteria. For MR11, which involves numerical scores ranging from 1 to 5, a metamorphic test is marked as biased when the difference between source and follow-up responses equals or exceeds 3. MR12 and MR13 use answer equivalence, flagging cases as biased if responses differ (i.e., case-insensitive exact match). For MR14, which compares ranked lists, we employed Spearman rank correlation, with values below 0.3 indicating weak agreement [54], and thus bias.
@@
MR1 - MR7 LLM-as-a-judge judgement(P,P,R,R)($P^{\prime},$$R^{\prime}$)
[…]
@@ 4 Tooling
Our approach is supported by an ecosystem of three tools: MUSE [14], GUARD-ME [16], and GENIE [15] (Figure 1). Together, these tools make our methodology actionable by both researchers and practitioners. They are implemented in TypeScript, comprising a codebase of over 5K lines. Each tool supports multiple deployment modalities, including direct execution via Node.js or containerised deployment using Docker. To facilitate adoption, the tools are distributed as REST APIs and come with comprehensive OpenAPI documentation and Postman collections.
@@ 5.1 Experimental Data
In early experiments, we used LLMs to generate a diverse range of demographic attributes based on the bias dimensions under study. However, we found that models like GPT-4o tended to overrepresent certain attributes, such as non-binary for gender or bisexual for sexual orientation, resulting in limited overall diversity. To address this issue, we manually created a catalogue of attributes to ensure broader and more representative metamorphic tests. The final catalogue comprises 190 attribute values. Of these, 120 correspond to demographic attributes spanning the five bias dimensions, while the remaining 70 are proper nouns associated with two genders and five religions. To maximise diversity, we sourced attributes from multiple references: for gender and sexual orientation, we used the LGBTQIA++ Wiki [56, 57], and for religion, socioeconomic status, and physical appearance, we relied on datasets compiled by Wan et al. [44]. For MR5, which relies on proper nouns, we compiled a diverse set of names from two sources: 1) the 20 most common male and female names in the United States over the past century [58], and 2) the 10 most popular names associated with each of the five major religions [59].
@@ 5.2.1 Experimental Setup
We selected a diverse set of LLMs to act as judges, focusing on widely adopted models of various sizes and including both open-source and commercial options. Our selection spanned five parameter ranges: 3-10B, 10-20B, 20-50B, 50-100B, and over 100B. Within the 3-10B range, we included Mistral and Llama 3.1, with 7B and 8B parameters, respectively. The 10-20B range featured Gemma 3 and DeepSeek-R1 (a distilled version from Qwen2.5), with 12B and 14B parameters, respectively. For the 20-50B range, we selected QwQ and DeepSeek-R1 (also distilled from Qwen2.5), both with 32B parameters. Although a 70B variant of DeepSeek-R1 is available, we opted for the 32B version because the larger model is a distilled version of Llama 3.3, which was already included in the 50-100B range. In the range of more than 100B, we included Gemini 2.0 Flash Thinking and o3-mini. Finally, we also investigated whether aggregating the decisions from three models, among those previously mentioned, could help smooth out inconsistencies and provide a more balanced and stable evaluation outcome.
@@ 5.2.2 Experimental Results
Table 3 presents the effectiveness of individual judge models and the top five (out of 84) evaluated three-model groups based on F1-score in bias detection. For each model or group of models, we report precision, recall, F1-score, and ROC AUC, using weighted averages to account for the class imbalance between biased and unbiased cases. We consider biased responses correctly identified as biased by the LLM as true positives (TP), unbiased responses correctly identified as unbiased as true negatives (TN), unbiased responses incorrectly reported as biased as false positives (FP), and biased responses incorrectly reported as unbiased as false negatives (FN). Thus, precision measures the proportion of responses flagged as biased that are truly biased, reflecting the reliability of the bias detection capabilities of the model (higher values mean fewer false alarms). Recall measures the proportion of all truly biased responses that are successfully identified, reflecting the ability of the model to capture most instances of bias (higher values mean fewer missed cases). In addition, we introduce a stability metric, defined as the proportion of cases in which a model produced consistent verdicts across the three independent executions.
The highest F1-scores across model size categories were obtained by Mistral 7B (0.73) in the 3-10B range, Gemma 3 12B (0.70) in the 10-20B range, R1-Qwen 32B (0.71) in the 20-50B range, Llama 3.3 70B (0.77) in the 50-100B range, and Gemini 2.0 Flash Thinking (0.74) in the 100B+ range. Llama 3.3 (70B) achieved the highest overall individual performance with an F1-score of 0.77 and an AUC of 0.73. At the MR level, its F1-scores ranged from 0.58 to 0.96, with the highest values observed for MR6 and the lowest for MR8. In contrast, performance across bias dimensions was more consistent, with F1-scores ranging from 0.66 to 0.87. Biases related to physical appearance were detected most reliably, whereas those associated with sexual orientation proved more challenging.
Notably, smaller models such as Mistral (7B) delivered competitive results, outperforming several models with a significantly larger number of parameters. In terms of stability, most models demonstrated high consistency (>0.9), with Gemini 2.0 Flash Thinking showing the highest score (1.00). Exceptions were observed in the R1-Qwen variants of 14B and 32B, which exhibited lower stability scores of 0.75 and 0.71, respectively.
@@
Table 4 summarises a statistical comparison between Llama 3.3 (70B), the best-performing individual judge model, and the remaining models. We used chi-squared tests with Bonferroni correction to evaluate differences and reported Cramér’s V as the effect-size measure. The results show that Llama 3.3 (70B) significantly outperformed R1-Qwen 32B, R1-Qwen 14B, and Gemini 2.0 Flash Thinking overall. Specifically, it achieved significantly higher scores than R1-Qwen 32B on MR2 - MR6, and than R1-Qwen 14B on MR2, MR4, and MR6. In addition, Llama 3.3 (70B) showed significantly superior performance compared to both Llama 3.1 (8B) and Gemma 3 (12B) on MR1, and to OpenAI o3-mini on MR2. The complete set of pairwise comparisons across all model combinations, along with a detailed per-MR analysis for the best-performing judge, is provided in the supplementary material [17].
Table 4: Statistical comparison between Llama 3.3 (70B) against the opponent judge models. Asterisks (*) indicate statistically significant differences.
Opponent model p-value p-adjusted Cramér’s V
R1-Qwen 32B 1.02e-21 8.18e-21* 0.31
R1-Qwen 14B 1.31e-16 1.05e-15* 0.27
Gemini 2.0 Flash Thinking 8.20e-06 6.56e-05* 0.15
Mistral (7B) 6.26e-03 5.01e-02 0.09
Gemma 3 (12B) 7.10e-03 5.68e-02 0.09
Llama 3.1 (8B) 3.38e-01 1.00e+00 0.03
OpenAI o3-mini 4.04e-01 1.00e+00 0.03
QwQ (32B) 4.83e-01 1.00e+00 0.02